OpenStudy (anonymous):

The length of a rectangle is 3 yd more than twice its width, and the area of the rectangle is 35 yd^2 Find the dimensions of the rectangle.

OpenStudy (anonymous):

A = w * l = w * ( 3 + 2w) = 35 yd² Can you continue

OpenStudy (anonymous):

lemme try

OpenStudy (anonymous):

Post your solution, I'll be back to check it :)

OpenStudy (anonymous):

this is what came up with me 3w+2w^2-35=0

OpenStudy (anonymous):

Great, can you solve for w? ( just like x)

OpenStudy (anonymous):

you want me to factor it ?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

ok i got it

OpenStudy (anonymous):

-5,7/2

OpenStudy (anonymous):

Wow, you're fast :D

OpenStudy (anonymous):

where is the width and the length ?

OpenStudy (anonymous):

Just take positive value width w = 3.5 ft Can you plug it back to find length?

OpenStudy (anonymous):

how did you get 3.5 ?

OpenStudy (anonymous):

7/2

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so the equation its fonna be like this : 3.5=35.L

OpenStudy (anonymous):

Check my first post to find out L = ...?

OpenStudy (anonymous):

L=A/W

OpenStudy (anonymous):

Nice try, but look more carefully :P

OpenStudy (anonymous):

l=w/a ?

OpenStudy (anonymous):

My first post, the second line! What do you understand?

OpenStudy (anonymous):

You already had the width, now find out the relationship of width and length to => L

OpenStudy (anonymous):

i cant get it

OpenStudy (anonymous):

"The length of a rectangle is 3 yd more than twice its width" => L = 3 + 2w

OpenStudy (anonymous):

so the length is 5 ?

OpenStudy (anonymous):

What is your width?

OpenStudy (anonymous):

3.5?

OpenStudy (anonymous):

Length L = 3 + 2 * 3.5 = ....